OpenStudy (anonymous):

If 2 over 5 of a ton of concrete covers 5 over 6 of a wall, how many tons of concrete are required to cover the entire wall? 12 over 25 ton 3 over 5 ton 1 and 7 over 30 tons 2 and 1 over 12 tons

OpenStudy (anonymous):

help pls

OpenStudy (acxbox22):

2/5 x ---- = ------ 5/6 1 2/5=5/6x x=12/25 tons

OpenStudy (anonymous):

If a sprinkler waters 1 over 14 of a lawn in 1 over 2 hour, how much time will it take to water the entire lawn? 1 over 7 hour 4 over 7 hour 4 hours 7 hours

OpenStudy (acxbox22):

post each question in a new thread please